# Average entropy of Gaussian mixtures

### v1

@@ Abstract
We calculate the average differential entropy of a $q$-component Gaussian mixture in $\mathbb{R}^{n}$. For simplicity, all components have covariance matrix $\sigma^{2}{\mathbf{1}}$, while the means $\{\mathbf{W}_{i}\}_{i=1}^{q}$ are i.i.d. Gaussian vectors with zero mean and covariance $s^{2}{\mathbf{1}}$. We obtain a series expansion in $\mu=s^{2}/\sigma^{2}$ for the average differential entropy up to order $\mathcal{O}(\mu^{2})$, and we provide a recipe to calculate higher order terms. Our result provides an analytic approximation with a quantifiable order of magnitude for the error, which is not achieved in previous literature.
@@ 1 Introduction
1.1 Entropy of Gaussian mixture distributions
A Gaussian mixture probability density on $\mathbb{R}^{n}$ is a function of the form $f(x)=\sum_{i=1}^{q}p_{i}g_{w_{i},K_{i}}(x)$, where $q$ is some integer, the $p_{i}$ are probabilities, and $g_{w,K}$ stands for the Gaussian distribution $g_{w,K}(x)=(2\pi)^{-\frac{n}{2}}(\det K)^{-\frac{1}{2}}\exp[-\frac{1}{2}(x-w)^{T}K^{-1}(x-w)]$, with $x,w\in\mathbb{R}^{n}$. In words, the density $f(x)$ is built as a weighted average of $q$ different $n$-dimensional Gaussian distributions. A mixture like this occurs when one stochastic process, with probability mass function $p_{1},\ldots,p_{q}$ determines which distribution is chosen for $x$. Analytically computing or estimating the differential entropy of a Gaussian mixture is a difficult problem. The differential entropy of a continuous variable $X\in\cal X$, $X\sim f$ is defined as $h(X)=-\int_{\cal X}f(x)\ln f(x){\rm d}x$ [1]. In the special case of a single Gaussian component ($q=1$), the integral can be computed analytically, yielding $\frac{n}{2}\ln 2\pi e+\frac{1}{2}\ln\det K$. In [6], the authors study the case of $q=2$, $n=1$, and $w_{1}=-w_{2}$, where a numerical approximation for this case is obtained. It is a known fact that given second moment of a random variable, the distribution with maximum entropy is the Gaussian distribution [1]. That is, we can easily obtain a loose upper bound for the differential entropy of a Gaussian mixture given by [8]:
@@ 1.1 Entropy of Gaussian mixture distributions
where $\Sigma$ is given by:
@@
One can then find approximations to the relative entropy as in [3, 2]. Although this method is efficient, it also does not a have closed-form expression and only provides an upper bound to $h(X)$. Other Monte Carlo sampling methods give guaranteed convergence. However, they become computationally demanding as the number of samples grows. An approximation for $h(X)$ was obtained in [4] by performing a Taylor series expansion of the logarithm term. In order to avoid the need to include higher order terms, a splitting of Gaussian components belonging to the density outside the logarithm is applied. The idea is to split components with high variance and replace them with Gaussian mixtures with components of lower variance. This is because the higher the variance of the components the higher the number of terms needed in the Taylor expansion to achieve a certain level of accuracy. Of course, this process is not exact and produces an error depending on how many Gaussian components are included in the split. The splitting method is not directly comparable to our work as we do not make a distinction between variances of the different components in the mixture, instead our expansion parameter is how spread the components of the mixture are relative to their shared variance value. However, one can obtain a basic upper bound on $h(X)$ as shown in [4]:
@@
As mentioned in [4], the upper bound (4) is significantly tighter than the Gaussian bound (1); it is exact in the case of a single Gaussian; and it is arbitrarily close to the real value of $h(X)$ when the support shared between the components of the mixture is negligible. A refinement of this bound was also introduced by means of merging different clusters in the mixture. However, we are concerned with the average over all possible mixtures, and therefore this family of bounds is less comparable to our result. In [5] the differential entropy is estimated for mixture distributions using pair-wise distances between the components of the mixture. It is shown that the estimator is a lower bound when the Chernoff $\alpha$-divergence is used, and an upper bound when the Kullback–Leibler divergence is used. From the results of [5], it is difficult to obtain tight bounds on the average entropy, which is the focus of our paper.
@@ 1.2 Contributions
We develop an analytical estimation method for the average Gaussian mixture entropy problem in the special case of equal weights $\frac{1}{q}$ and equal covariance matrix $\sigma^{2}{\mathbf{1}}$. Our method postulates that the displacements $w_{i}$ themselves are stochastic, i.i.d. with Gaussian distribution $g_{0,s^{2}{\bf 1}}$. We compute $h(X)$ averaged over the displacements $w_{i}$. We work in the regime $s<\sigma$. Our method uses the fraction $\mu=\frac{s^{2}}{\sigma^{2}}$ as the small parameter for a power expansion. We show results up to and including order $\mathcal{O}(\mu^{2})$. Our result is novel since most available estimators of $h(X)$ in the literature have no closed-form expressions, and usually rely on upper bounds rather than an explicit calculation of $h(X)$. In [4] although a Taylor series approximation is employed, its accuracy is dependent on the splitting mentioned in Section 1.1. Furthermore, it is difficult to estimate the order of magnitude for the error in a general expansion around the means. In contrast, our setup does not need the splitting method, and for our result it is possible to quantify the order of magnitude for the error since our power series is in $\mu$.
@@ 2.1 Notation
Boldface lowercase letters denote $n$-dimensional vectors (e.g. $\mathbf{a}$); the inner product between two $n$-dimensional vectors $\mathbf{a}$ and $\mathbf{b}$ is denoted by $\mathbf{a}\cdot\mathbf{b}$; A hat denotes an array in $q$ dimensions, i.e. indexed with the number of the Gaussian components; for lowercase letters $\hat{(\cdot)}$ denotes a $q$-dimensional vector (e.g. $\hat{a}$); for uppercase letters a $q\times q$ matrix (e.g. $\hat{A}$). $\hat{a}^{\mathrm{T}}\hat{b}$ denotes $q$-dimensional inner product. $\otimes$ denotes the Kronecker product.
@@ 2.2 Starting point
We note that a simplification is possible when $n>q$, i.e. the dimension of the space is higher than the number of components in the mixture. For all configurations $\hat{\mathbf{w}}$ it is possible to find a rotation such that after rotation ${\mathbf{w}}_{j}=\sum_{\alpha=1}^{q}w_{j\alpha}{\mathbf{e}}_{\alpha}$, where the ${\mathbf{e}}_{\alpha}$ are the basis vectors in ${\mathbb{R}}^{n}$. In other words, for all ${\mathbf{w}}_{j}$ the vector components in the dimensions beyond $q$ vanish. The density (5) simplifies to
@@
The first exponent is trivially dealt with when the $\ln f_{{\mathbf{X}}|{\hat{\mathbf{W}}}}({\mathbf{x}}|\hat{\mathbf{w}})$ is integrated, yielding a constant contribution $\frac{n-q}{2}$ to the entropy. Hence we will focus only on the case $n\leq q$ in our calculations. In general, analytically approximating the differential entropy $h({\mathbf{X}}|\hat{\mathbf{W}}=\hat{\mathbf{w}})$ for given $\hat{\mathbf{w}}$ is a difficult problem. In this paper we will study a problem that is slightly easier: finding analytic approximations for $h({\mathbf{X}}|\hat{\mathbf{W}})={\mathbb{E}}_{\hat{\mathbf{w}}}h({\mathbf{X}}|\hat{\mathbf{W}}=\hat{\mathbf{w}})$ in the case where the offsets ${\mathbf{w}}_{j}$ are i.i.d. Gaussian-distributed with zero mean and covariance matrix $s^{2}{\bf 1}$.
@@
We will take $s^{2}<\sigma^{2}$ and develop a power series in the parameter $\mu=s^{2}/\sigma^{2}$. In the following sections we obtain our result by following these steps:
  • We first obtain a more compact form of $h({\mathbf{X}}|\hat{\mathbf{W}})$ in Corollary 2.3 via a change a variables.
  • After performing another change of variables to remove correlations, we evaluate the leading terms to $h({\mathbf{X}}|\hat{\mathbf{W}})$.
  • In Theorem 2.7 we obtain an expression that separates the leading contributions to $h({\mathbf{X}}|\hat{\mathbf{W}})$ and the quantity $S$ to be defined.
  • We perform a third change of variables to simplify $S$, and we obtain a Taylor series for $h({\mathbf{X}}|\hat{\mathbf{W}})$ up to order $\mathcal{O}(\mu^{2})$ in Theorem 3.1 via a brute-force approach.
  • Finally, we provide a determinant based approach to evaluate the power series for $S$, and we obtain a result for $h({\mathbf{X}}|\hat{\mathbf{W}})$ up to order $\mathcal{O}(\mu)$.
@@ 2.3 First change of variables
The differential entropy $h({\mathbf{X}}|\hat{\mathbf{W}})$ is given by:
@@
Next we can get rid of the $\sum_{j}$ summation, using permutation symmetry.
@@
Let $F(\mathbf{z}_{1},\dots,\mathbf{z}_{q})$ be a any permutation invariant function, then we have:
@@
where the matrix elements $\{M^{(k)}_{ij}\}_{k=1}^{q}$ are given by:
@@
then the differential entropy $h({\mathbf{X}}|\hat{\mathbf{W}})$ can be written as:
@@
It follows directly from Lemma 2.1 and Lemma 2.2, where we note that equation (11) is permutation invariant. ∎
@@ 2.4 Diagonalization of $\hat{M}^{(q)}$ and second change of variables
We would like to obtain an expression for the expectation term in equation (12), however, we first need to switch to a diagonal form that removes correlation between the integration variables, i.e. we would like to find the eigenvectors and eigenvalues of the matrix $\hat{M}^{(q)}$. We first notice that adding the identity matrix, while it shifts the eigenvalues by 1, does not change the eigenvectors of a matrix. Hence we diagonalize the matrix $\hat{C}^{(q)}$ as shown in Theorem 2.4, and our result follows directly in Corollary 2.5.
@@
Let $\hat{J}$ be the all-one matrix. Let $\hat{e}^{(k)}$ denote the $k$-th standard basis vector. Then the matrix $\hat{C}^{(q)}$ given by:
@@
has an orthogonal matrix $\hat{\Lambda}^{(q)}$ of eigenvectors given by:
@@
where $\{\hat{\alpha}_{i}\}^{q-2}_{i=1}$ are eigenvectors with eigenvalue 0, whereas $\hat{\beta}_{1}$ and $\hat{\beta}_{2}$ have eigenvalues $\lambda_{1}$ and $\lambda_{2}$, respectively, and read:
@@
Furthermore,the elements of the eigenvectors are given by:
@@
Note that in matrix notation, we can write:
@@
where $\hat{I}$ denotes the identity matrix, and we have the following corollary.
@@
The matrix $\hat{M}^{(q)}$ given by equation (18) is diagonalizable by the matrix of eigenvectors $\hat{\Lambda}^{(q)}$ given in Theorem 2.4, and we have:
@@
Let $\hat{\mathbf{z}}=\hat{\Lambda}^{(q)}\hat{\mathbf{u}}$, then the function $F$ given in Corollary 2.3 can be written as:

### v2

@@ Abstract
We calculate the average differential entropy of a $q$-component Gaussian mixture in $\mathbb{R}^{n}$. For simplicity, all components have covariance matrix $\sigma^{2}{\mathbf{1}}$, while the means $\{\mathbf{W}_{i}\}_{i=1}^{q}$ are i.i.d. Gaussian vectors with zero mean and covariance $s^{2}{\mathbf{1}}$. We obtain a series expansion in $\mu=s^{2}/\sigma^{2}$ for the average differential entropy up to order $\mathcal{O}(\mu^{2})$, and we provide a recipe to calculate higher-order terms. Our result provides an analytic approximation with a quantifiable order of magnitude for the error, which is not achieved in previous literature.
@@ 1 Introduction
1.1 Gaussian Mixtures
A Gaussian mixture probability density on $\mathbb{R}^{n}$ is a function of the following form:
$f(x)=\sum_{i=1}^{q}p_{i}g_{w_{i},K_{i}}(x),$ (1)
where $q$ is some integer, $p_{i}$ are probabilities, and $g_{w,K}$ stands for the Gaussian distribution $g_{w,K}(x)=(2\pi)^{-\frac{n}{2}}(\det K)^{-\frac{1}{2}}\exp[-\frac{1}{2}(x-w)^{T}K^{-1}(x-w)]$, with $x,w\in\mathbb{R}^{n}$. In words, the density $f(x)$ is built as a weighted average of $q$ different $n$-dimensional Gaussian distributions. A mixture like this occurs when one stochastic process, with probability mass function $p_{1},\ldots,p_{q}$, determines which distribution is chosen for $x$. Gaussian mixtures are widely used in various areas for their simplicity, as well as their wide range of applicability by virtue of being global (smooth) function approximators. Most recently, in astrophysics, Gaussian mixture models were used to study the kinematics of dark matter [23], as well as predicting the spectrum of quasars [20]. Furthermore, Gaussian mixtures are also widely used in the intersection between statistical physics and machine learning, i.e., diffusion models [22, 6]. Diffusion models became popular for their wide range of applications, including generative tasks and image restoration, for example, see [17, 4] for Gaussian mixture diffusion models. In wireless communications, Gaussian mixtures were recently utilized to estimate the channel trajectories of moving mobile terminals [19], while in cybernetic security, they played a role in defending against the Byzantine attack [13], and earlier for authentication in wireless networks [14]. In bioinformatics, they are used as an alternative to clustering algorithms for analyzing gene expression data [9, 18]. They also find application in thermodynamics, e.g., a computationally efficient Monte Carlo method was devised in [15] to study the properties of nonadiabatic systems. In [16], many of the above-mentioned disciplines are combined, where the authors evaluate probabilities in deep generative models.
1.2 Related Work
[…]
@@ 1.2 Related Work
where $\Sigma$ is given by
@@
One can then find approximations to the relative entropy, as in [5, 3]. Although this method is efficient, it also does not a have closed-form expression and only provides an upper bound to $h(X)$. Other Monte Carlo sampling methods give guaranteed convergence. However, they become computationally demanding as the number of samples grows. An approximation for $h(X)$ was obtained in [7] by performing a Taylor series expansion of the logarithm term. In order to avoid the need to include higher-order terms, a splitting of Gaussian components belonging to the density outside the logarithm is applied. The idea is to split components with high variance and replace them with Gaussian mixtures with components of lower variance. This is because the higher the variance of the components the higher the number of terms needed in the Taylor expansion to achieve a certain level of accuracy. Of course, this process is not exact and produces an error depending on how many Gaussian components are included in the split. The splitting method is not directly comparable with our work, as we do not make a distinction between variances of the different components in the mixture; instead, our expansion parameter is how spread the components of the mixture are relative to their shared variance value. However, one can obtain a basic upper bound on $h(X)$, as shown in [7]:
@@
As mentioned in [7], the upper bound (5) is significantly tighter than the Gaussian bound (2); it is exact in the case of a single Gaussian; and it is arbitrarily close to the real value of $h(X)$ when the support shared between the components of the mixture is negligible. A refinement of this bound was also introduced by means of merging different clusters in the mixture. However, we are concerned with the average over all possible mixtures, and therefore, this family of bounds is less comparable with our result. In [8], the differential entropy is estimated for mixture distributions using pair-wise distances between the components of the mixture. It is shown that the estimator is a lower bound when the Chernoff $\alpha$-divergence is used, and an upper bound when the Kullback–Leibler divergence is used. From the results of [8], it is difficult to obtain tight bounds on the average entropy, which is the focus of our paper.
@@ 1.3 Contributions
We develop an analytical estimation method for the average Gaussian mixture entropy problem in the special case of equal weights $\frac{1}{q}$ and equal covariance matrix $\sigma^{2}{\mathbf{1}}$. Our method postulates that the displacements $w_{i}$ themselves are stochastic, i.i.d. with Gaussian distribution $g_{0,s^{2}{\bf 1}}$. We compute $h(X)$ averaged over the displacements $w_{i}$, resulting in the conditional entropy $h(X|W)$. We mention that the quantity $h(X|W)$, for exactly this setting with equal weights, spherical covariance, and Gaussian displacements, plays a role in the detectability of digital watermarks (see Section 2.2).
We work in the regime $s<\sigma$. Our method uses the fraction $\mu=\frac{s^{2}}{\sigma^{2}}$ as the small parameter for a power expansion. We show results up to and including order $\mathcal{O}(\mu^{2})$. Our result is novel since most available estimators of $h(X)$ in the literature have no closed-form expressions, and usually rely on upper bounds rather than an explicit calculation of $h(X)$. In [7], although a Taylor series approximation is employed, its accuracy is dependent on the splitting mentioned in Section 1.2. Furthermore, it is difficult to estimate the order of magnitude for the error in a general expansion around the means. In contrast, our set-up does not need the splitting method, and for our result, it is possible to quantify the order of magnitude for the error since our power series is in $\mu$.
@@ 2.1 Notation
Boldface lowercase letters denote $n$-dimensional vectors (e.g., $\mathbf{a}$); the inner product between two $n$-dimensional vectors $\mathbf{a}$ and $\mathbf{b}$ is denoted by $\mathbf{a}\cdot\mathbf{b}$; A hat denotes an array in $q$ dimensions, i.e., indexed with the number of the Gaussian components; for lowercase letters, $\hat{(\cdot)}$ denotes a $q$-dimensional vector (e.g., $\hat{a}$); for uppercase letters, a $q\times q$ matrix (e.g., $\hat{A}$). $\hat{a}^{\mathrm{T}}\hat{b}$ denotes the $q$-dimensional inner product. $\otimes$ denotes the Kronecker product.
@@ 2.2 Starting Point
We note that a simplification is possible when $n>q$, i.e., the dimension of the space is higher than the number of components in the mixture. For all configurations $\hat{\mathbf{w}}$, it is possible to find a rotation such that after rotation ${\mathbf{w}}_{j}=\sum_{\alpha=1}^{q}w_{j\alpha}{\mathbf{e}}_{\alpha}$, where the ${\mathbf{e}}_{\alpha}$ are the basis vectors in ${\mathbb{R}}^{n}$. In other words, for all ${\mathbf{w}}_{j}$, the vector components in the dimensions beyond $q$ vanish. The density (6) simplifies to
@@
The first exponent is trivially dealt with when the $\ln f_{{\mathbf{X}}|{\hat{\mathbf{W}}}}({\mathbf{x}}|\hat{\mathbf{w}})$ is integrated, yielding a constant contribution $\frac{n-q}{2}$ to the entropy. Hence, we only focus on the case $n\leq q$ in our calculations. In general, analytically approximating the differential entropy $h({\mathbf{X}}|\hat{\mathbf{W}}=\hat{\mathbf{w}})$ for a given $\hat{\mathbf{w}}$ is a difficult problem. In this paper, we study a problem that is slightly easier: finding analytic approximations for $h({\mathbf{X}}|\hat{\mathbf{W}})={\mathbb{E}}_{\hat{\mathbf{w}}}h({\mathbf{X}}|\hat{\mathbf{W}}=\hat{\mathbf{w}})$ in the case where the offsets ${\mathbf{w}}_{j}$ are i.i.d. Gaussian-distributed with zero mean and covariance matrix $s^{2}{\bf 1}$.
@@
We note that the quantity $h({\mathbf{X}}|\hat{\mathbf{W}})$, under precisely these conditions, occurs in the field of spread-spectrum watermarking [2, 21]. In this setting, the $\hat{\mathbf{W}}$ is a table of $q$ random watermarking sequences. A row index $K\in\{1,\ldots,q\}$ is chosen at random, and the $K$’th row of $\hat{\mathbf{W}}$ is additively inserted into some data ${\mathbf{D}}$, which are modeled as Gaussian, and then attackers apply noise in order to wipe out the watermark. The ${\mathbf{X}}$ plays the role of the watermarked data after this attack. A watermark detector tries to determine the index $K$ based on ${\mathbf{X}}$ and $\hat{\mathbf{W}}$ (and optionally the unwatermarked original data ${\mathbf{D}}$). In the analysis of the detection efficiency, an important figure of merit is the mutual information $I(K;{\mathbf{X}}\hat{\mathbf{W}})$ or $I(K;{\mathbf{D}}{\mathbf{X}}\hat{\mathbf{W}})$, both of which after some rewriting involve precisely the conditional entropy $h({\mathbf{X}}|\hat{\mathbf{W}})$. We take $s^{2}<\sigma^{2}$ and develop a power series in the parameter $\mu=s^{2}/\sigma^{2}$. In the following sections, we obtain our results by following these steps:
  • We first obtain a more compact form of $h({\mathbf{X}}|\hat{\mathbf{W}})$ in Corollary 2.3 via a change of variables in order to write $h({\mathbf{X}}|\hat{\mathbf{W}})$ in terms of the more familiar expectation with regard to a multivariate Gaussian density.
  • We next perform another change or variables to obtain a diagonal form in the expectation with regard to the multivariate Gaussian. This is important since our new variables will be independent, and many expressions will be simplified if our variables are not correlated.
  • We evaluate the leading terms to $h({\mathbf{X}}|\hat{\mathbf{W}})$, as shown in Theorem 2.7, where we obtain an expression that separates the leading contributions to $h({\mathbf{X}}|\hat{\mathbf{W}})$ and the quantity $S$ to be defined. The reason this is important is that the leading contributions contain terms that will make the Taylor series diverge, and therefore, we evaluate them analytically before performing the Taylor series expansion. We also include other terms that do not cause any problems in the limit of small $\mu$ for convenience. The remaining expression $S$ is then safe to expand for small $\mu$.
  • We perform a third change of variables to simplify $S$, and we obtain a Taylor series for $h({\mathbf{X}}|\hat{\mathbf{W}})$ up to order $\mathcal{O}(\mu^{2})$ in Theorem 3.1 via a brute-force approach. The change of variables is necessary for making the analytical expression tractable.
  • Finally, we provide a determinant-based approach to evaluate the power series for $S$, and we obtain a result for $h({\mathbf{X}}|\hat{\mathbf{W}})$ up to order $\mathcal{O}(\mu)$.
@@ 2.3 First Change of Variables
The differential entropy $h({\mathbf{X}}|\hat{\mathbf{W}})$ is given by
@@
Next, we can get rid of the $\sum_{j}$ summation, using permutation symmetry.
@@
Let $F(\mathbf{z}_{1},\dots,\mathbf{z}_{q})$ be any permutation invariant function; then, we have
@@
where the matrix elements $\{M^{(k)}_{ij}\}_{k=1}^{q}$ are given by
@@
then, the differential entropy $h({\mathbf{X}}|\hat{\mathbf{W}})$ can be written as
@@
It follows directly from Lemmas 2.1 and 2.2, where we note that Equation (12) is permutation invariant. ∎
@@ 2.4 Diagonalization of $\hat{M}^{(q)}$ and Second Change of Variables
We would like to obtain an expression for the expectation term in Equation (13); however, we first need to switch to a diagonal form that removes correlation between the integration variables, i.e., we would like to find the eigenvectors and eigenvalues of the matrix $\hat{M}^{(q)}$. We first notice that adding the identity matrix, while it shifts the eigenvalues by 1, does not change the eigenvectors of a matrix. Hence, we diagonalize the matrix $\hat{C}^{(q)}$, as shown in Theorem 2.4, and our result follows directly in Corollary 2.5.
@@
Let $\hat{J}$ be the all-one matrix. Let $\hat{e}^{(k)}$ denote the $k$-th standard basis vector. Then, the matrix $\hat{C}^{(q)}$ given by
@@
has an orthogonal matrix $\hat{\Lambda}^{(q)}$ of eigenvectors given by
@@
where $\{\hat{\alpha}_{i}\}^{q-2}_{i=1}$ are eigenvectors with eigenvalue 0, whereas $\hat{\beta}_{1}$ and $\hat{\beta}_{2}$ have eigenvalues $\lambda_{1}$ and $\lambda_{2}$, respectively, and read
@@
Furthermore, the elements of the eigenvectors are given by
@@
Note that in matrix notation, we can write
@@
where $\hat{I}$ denotes the identity matrix, and we have the following corollary:
@@
The matrix $\hat{M}^{(q)}$ given by Equation (19) is diagonalizable by the matrix of eigenvectors $\hat{\Lambda}^{(q)}$ given in Theorem 2.4, and we have
@@
Let $\hat{\mathbf{z}}=\hat{\Lambda}^{(q)}\hat{\mathbf{u}}$, then the function $F$ given in Corollary 2.3 can be written as